# Generative Visual Instruction Tuning

### v1

@@ Abstract
We propose to use machine-generated instruction-following data to improve the zero-shot capabilities of a large multimodal model with additional support for generative and image editing tasks. We achieve this by curating a new multimodal instruction-following set using GPT-4V and existing datasets for image generation and editing. Using this instruction set and the existing LLaVA-Finetune instruction set for visual understanding tasks, we produce GenLLaVA, a Generative Large Language, and Visual Assistant. GenLLaVA is built through a strategy that combines three types of large pretrained models through instruction finetuning: LLama for language modeling, SigLIP for image-text matching, and StableDiffusion for text-to-image generation. Our model demonstrates visual understanding capabilities on par with LLaVA and additionally demonstrates competitive results with native multimodal models such as Unified-IO 2, paving the way for building advanced general-purpose visual assistants by effectively re-using existing multimodal models. We open-source our dataset, codebase, and model checkpoints to foster further research and application in this domain [^1] https://github.com/jeffhernandez1995/GenLlaVA.git.
@@ 1 Introduction
Large multimodal models have become increasingly popular in the research community as they are one of the key building blocks for general-purpose assistants [1, 40, 2]. The recently proposed LLaVA [25] model is among the latest wave of works that have demonstrated the effectiveness of instruction tuning for multimodal models [23, 51, 41, 14]. There has been a considerable amount of work building on top of the LLaVA model ranging from image generation [17, 36, 35], grounding [44], image editing [7] to video understanding [20]. These works share the same principles; they extend the ideas of visual instruction tuning to one or more capabilities. However, after adding a new capability (i.e., image generation), the resulting models often lose some, if not most, of their visual and language understanding capabilities. This contrasts with the original visual instruction tuning in which the models retained their language capabilities after training for visual understanding.
@@
Figure 1: Comparison of GenLLaVA compared to recent architectures, unlike BLIP-2 [19], we use a Linear projector similar to the LlaVA architecture [25]. Generation capabilities are added using a diffusion model, but unlike GILL [17], we use a Q-former as the generation head. Finally, our model benefits from using a stronger visual encoder, namely SigLIP[49]; a stronger LLM, namely Mistral-7b [12]; and a stronger diffuser, namely SDv1.4 [33]. L stands for Linear projection, and Q stands for Q-former resampler.
@@
In this paper, we present the generative visual instruction tunining, in which we teach a Large Multimodal Model (LMM) image understanding, image generation, and image editing tasks without significantly diminishing the performance of each individual capability. (See Fig. 1 for an overview of our method). To our knowledge, this is the first time such capability has been achieved, and our findings pave the way for building a general-purpose visual assistant. Our contributions are the following:
@@
  • Generative multimodal instruction-following data. Inspired by previous work [25], we curate a multimodal instruction tuning set that combines image understanding, image generation, and image editing data.
  • New improved training recipe. We empirically design a single-phase training recipe to teach the LMM the necessary capabilities, unlike its predecessor, which requires multi-phase training recipes.
@@
  • Open source. We release the following assets to the public: the generated multimodal instruction data, the codebase, the model checkpoints, and a visual chat demo.
@@ 2 Related Work
Large Multimodal Models (LMMs)
Large Multimodal Models (LMMs) refer to large language models that can understand various modalities beyond human language. Some research efforts are focused on combining image, audio, video, and other modalities with language  [50, 27], while others aim to enhance the fusion of vision knowledge and language. For example, BLIP-2 [19] created a large-scale image captioning dataset and paired a language model with a vision encoder to produce a robust multimodal model. Following this, LLaVA [25] developed a cost-effective approach to train an advanced LLM through visual instruction tuning. Although LLaVA-Next [24] improved performance for single-image tasks, it required over 2,000 tokens per image, which is about four times more than the original LLaVA. More recent models such as QwenVL[2], CogVLM[42], and Yi-VL[45] follow architectures similar to those of LLaVA. Our proposed method does not only focus on models for multimodal understanding, but also on adding generative capabilities to such models.
Generation with LMMs
In recent years, the design of methods that aim to adapt LLMs and LMMs for generative tasks has become an active research direction. For instance, GILL [17] translates the hidden representations of an LLM into embeddings that correspond to a text-to-image model by learning a neural network to perform efficient mapping using the text encoder of the diffusion model. AnyGPT [50] discretizes data from multiple modalities, extends the existing LLM vocabulary to add the extra modalities, and incorporates new randomly initialized parameters that enable additional input embeddings and prediction outputs. MGIE [7] adapts the text embedder, image input adapter, and LM head output parameters of an LMM jointly with a diffusion model for image editing from instructions. Codi-2 [38] follows a similar line and adds multiple decoders for image, audio, and video generation to condition an LLM to generate interleaved outputs from different modalities. CM3leon [46] proposes a token-based decoder-only LMM based on the CM3 architecture that is capable of both text and image generation and editing. Unified-IO 2 [27] and Chameleon [39] and GPT-4o take the early-fusion fully multimodal approach training the model from scratch to be able to expand the number of supported tasks and modalities.
@@ 3 Method
Figure 2: Editing capabilities of our model. GPT4 currently uses a version of the DALLE text-to-image model as a tool and hence is not directly able to edit images. GPT4o instead uses tools through Python generated code to accomplish the requested action. Our model, GenLLaVA, connects input features obtained from CLIP to a language model that also produces output embeddings for a text-to-image StableDiffusion model, achieving an end-to-end editing task with a multimodal model.
@@ 3.1 Background: Large Multimodal Models
Large language models (LLMs) excel in natural language generation. Large Multimodal Models enhance LLMs with the ability to interpret images and respond accordingly. Built upon a pre-trained LLM, the LMM incorporates a visual encoder (e.g., CLIP/L [31]) to derive visual features $f$, along with an adapter $\mathcal{W}$ that maps $f$ into the language domain. Following the training methodology of LLaVA [25], this process is encapsulated in the equation:
@@
where $l$ represents the length of the word tokens within $\mathcal{C}$. The set $\mathcal{C}$ can represent an image caption (Features Alignment) or multimodal instruction-following data (Instruction Tuning). The LMM employs the standard autoregressive method for next-token prediction, allowing it to function as a visual assistant across diverse tasks such as visual question answering and complex reasoning. Despite gaining visual perceptive abilities through this training, its responses are currently constrained to text.
@@ 3.2 Visual Generation in Large Multimodal Models
We append $N$ visual tokens [IMG] after the instruction, with their word embeddings being trainable. The LMM learns to generate these tokens through its language modeling (LM) head. These visual tokens represent visual-related instruction comprehension within $\mathcal{E}$ and form a bridge between the language and vision modalities. We follow the same visual generation framework of GILL [17] and MGIE [7] in extracting visual features, which we summarize here for succintness.
@@
Table 1: Main result. Comparison of various models across advanced knowledge and general understanding. MGIE was not originally designed for these tasks, and it is purely an editing model; we take the generated caption as the answer for VQA and give a blank image as a prompt when we ask it to generate completely new images. We intend to show that models lose previous capabilities when we add a new one.
Model name Advanced Knowlege General Understanting Editing Generation
MathVista MMMU MMVet SEED-B MMB EVR CC3M COCO
GILL [17] 18.6 28.8 23.3 52.5 38.2 30.4 15.3 0.67
AnyGPT [50] 24.4 30.6 33.1 44.5 36 40.3 14.3 0.65
MGIE [7] 15.5 25.6 13 28.8 6.6 71.5 13.6 0.66
Unified-IO 2 [27] 30.7 35 36.6 61.8 71.5 50.2 13.4 0.72
GenLLaVA (Ours) 24.9 29.7 33.1 63.5 65 64.7 14.3 0.71
To guide image generation with the visual latent information $\mathcal{U}$, we employ a latent diffusion model [32], incorporating a variational autoencoder (VAE) for handling denoising diffusion in the latent space. First, we encode the desired visual output via the diffusion model encoder $o=\text{Enc}_{\text{VAE}}(\mathcal{O})$; this output may be intended for image generation or editing tasks. The diffusion process progressively introduces noise into $o$ as $z_{t}$, increasing the noise level over timesteps $t$. We then train the UNet $\epsilon_{\theta}$ to predict the added noise [10]. The diffusion process is conditioned on the visual latent information $\mathcal{U}$ through a cross-attention layer, defined as $\text{Attention}(Q,K,V)=\text{softmax}(\frac{QK^{T}}{\sqrt{\text{dim}}})\cdot V$, where:
$Q=W^{(i)}_{Q}\cdot\varphi_{i}(z_{t}),K=W^{(i)}_{K}\cdot\{u\},V=W^{(i)}_{V}\cdot\{u\},$ (3)
with $\varphi$ representing the flattening operation, and $W^{(i)}_{Q}$, $W^{(i)}_{K}$, and $W^{(i)}_{V}$ being learnable attention matrices. We apply classifier-free guidance [11], where the score estimation $s_{\theta}$ is extrapolated to deviate from the unconditional $\varnothing$, following standard practices in diffusion models.
@@ 4 Experiment Settings
We use the ShareGPT [34] dataset. We filter out invalid conversations, but unlike Vicuna, long conversations that surpass 2048 tokens are truncated rather than split into multiple conversations. This results in $\sim$40K conversations
@@
For image generation, we use the original LLaVA-Pretrain, which we now invert to be an image generation dataset instead of an image caption dataset; this dataset contains $\sim$558K data points collected from the LAION, SBU, and CC3M datasets and captioned by the BLIP-2 model.
Image Editing. [9.63%]
We subsample the IPr2Pr [3] dataset as our editing data. We use $\sim$200K CLIP-filtered data, where instructions are extracted by GPT-3, and images are synthesized by the Prompt-to-Prompt model [8].
Image Understanting.[61.56%]
@@
We combine the original LLaVA-Finetune dataset with the LVIS-INSTRUCT4V[41], a new visual instruction tuning dataset constructed in the same way as the original LlaVA dataset but using GPT4-V as the captioner instead of BLIP-2. If they exist, we remove duplicates from the LLaVA-Finetune dataset; this results in $\sim$880K samples. We additionally add the following instruction datasets LRV-Instruction [22] ($\sim$80K), laion-gpt4v-dataset ($\sim$15K), ShareGPT4V[5] ($\sim$100K), and the following document understanding datasets DocVQA[30]($\sim$50K), SynDog-EN[16]($\sim$65K), ChartQA[29]($\sim$23K), DVQA[13] ($\sim$50K)and AI2D[15] ($\sim$15K).
4.1 Generative Visual Instruction Data
@@
4.2 Training details.
In this section, we evaluate our model on a broad range of tasks that require visual understanding and generation. We do not perform task-specific finetuning in any experiments. The Supplementary section details details about experimental setups, additional results, and additional studies on GenLLaVA ’s instruction capabilities.
We adopt LLaVA-v1.5-7B [23] architecture and pre-training weights, then tune it on the constructed GVIT-mix-2076K. We named this model GenLLaVA. Specifically, we replace LLaVA-v1.5-mix-665K with our GVIT-mix-2076K, add the visual generation module, and only perform a single-stage instruction tuning training as this was shown to be superior [14]. The rest of the model training protocol is kept unchanged for fair comparison. Generative Visual instruction tuning takes about 48 hours for both full-parameter tuning and LoRA tuning on 8 NVIDIA Tesla A100 GPUs, each with 80GB memory, with DeepSpeed ZeRO Stage 3.
Visual Understanting
We evaluate vision language performance and compare it against other generalist models, i.e., models that can do visual generation and understating. Results on a collection of 8 vision/language benchmarks are shown in Table 1. These benchmarks are designed to probe advanced knowledge and general understanding. MMBench [26] evaluates answer robustness through comprehensive shuffling of multiple-choice options. SEED-Bench [18] tests model performance on images and videos with multiple-choice questions. MM-Vet [47] evaluates the ability to engage in visual conversations and assess the accuracy and helpfulness of responses. Mathvista [28] consolidates mathematical reasoning benchmarks, focusing on logical and algebraic reasoning with puzzle tests. MMMU [48] covers 57 subjects across STEM, humanities, and social sciences, ranging from elementary to advanced professional levels, and tests both world knowledge and problem-solving skills.
@@
For visual generation evaluation, we measured the Fréchet Inception Distance (FID) [9] on the CC3M validation set [4] (which is a measure of image realism) and CLIP Similarity on the MS-COCO dataset [21] (which is a measure of the alignment between the text prompt and the generated image). For image editing, we measured the DINOScore on the 5.7K examples of the EVR dataset [37] following the same protocol of Fu et.al. [7].
Figure 4: Ablation on the effect of adding extra data to our instruction set.

### v2

@@ Abstract
We propose to use automatically generated instruction-following data to improve the zero-shot capabilities of a large multimodal model with additional support for generative and image editing tasks. We achieve this by curating a new multimodal instruction-following set using GPT-4V and existing datasets for image generation and editing. Using this instruction set and the existing LLaVA-Finetune instruction set for visual understanding tasks, we produce GenLLaVA, a Generative Large Language and Visual Assistant. GenLLaVA is built through a strategy that combines three types of large pretrained models through instruction finetuning: Mistral for language modeling, SigLIP for image-text matching, and StableDiffusion for text-to-image generation. Our model demonstrates visual understanding capabilities superior to LLaVA and additionally demonstrates competitive results with native multimodal models such as Unified-IO 2, paving the way for building advanced general-purpose visual assistants by effectively re-using existing multimodal models. We open-source our dataset, codebase, and model checkpoints to foster further research and application in this domain [^1] https://github.com/jeffhernandez1995/GenLlaVA.git.
@@ 1 Introduction
The field of multimodal models has become increasingly popular in the research community as they are one of the key building blocks for general-purpose assistants (Achiam et al., 2023; Gemini Team et al., 2023; Bai et al., 2023). One of the main directions researchers have pursued is to combine Large Language Models (LLMs) with Vision Models for multimodal tasks i.e. creating LVLMs. The recently proposed LLaVA model (Liu et al., 2023b) is among the latest wave of works that have demonstrated the effectiveness of instruction tuning for multimodal models (Liu et al., 2024a; Zhao et al., 2023; Wang et al., 2023a; Karamcheti et al., 2024). In these works, a two-stage pipeline is followed: (1) multimodal pre-training where the unimodal models are combined and trained on a large corpus of captioning data Schuhmann et al. (2022); Ordonez et al. (2011); and (2) a supervised fine-tuning (SFT) Liu et al. (2024a, b) stage where the model is trained on domain-specific data and enables it to better perform various downstream tasks of interest.
Visual generation is another research direction that combines visual and language modalities. There are two common approaches for text-to-image generation. One approach employs diffusion models (Rombach et al., 2022), which have shown unprecedented performance in image synthesis, becoming the de facto method for visual generation. The other line of work converts visual content into discrete tokens using vector quantization (VQ) and then leverages an autoregressive transformer for high-quality image synthesis (Chang et al., 2022; Lee et al., 2022).
As visual understanding and generation capabilities advance rapidly and independently, a growing trend is to combine these into a unified Large Multimodal Model (LMM). There are two main approaches to achieving such unification. Many LMMs (Koh et al., 2024; Sun et al., 2024b; Fu et al., 2024) produce conditional embeddings to be used by a pretrained diffusion model for image generation. On the other hand, there are LMMs (Lu et al., 2024a; Chameleon Team, 2024; Yu et al., 2023) that adopt VQ encoders to project visual inputs into discrete tokens and use the same next-token prediction paradigm as Language Models.
@@
There has been a considerable amount of work building on top of the LLaVA model ranging from image generation (Koh et al., 2024; Sun et al., 2024b, a), grounding (You et al., 2024), image editing (Fu et al., 2024) to video understanding (Lin et al., 2024). These works share the same principles; they extend the ideas of visual instruction tuning to one or more capabilities. However, after adding a new capability (i.e., image generation), the resulting models often lose some, if not most, of their visual and language understanding capabilities. We propose a strategy that leads to a model that can perform generative tasks while retaining multimodal understanding capabilities.
@@
Figure 1: Comparison of GenLLaVA against recent architectures. Unlike BLIP-2 (Li et al., 2023), we use a Linear projector similar to the LlaVA architecture (Liu et al., 2023b). Generation capabilities are added using a diffusion model, but unlike GILL (Koh et al., 2024), we use a Q-former as the generation head. Finally, our model benefits from using a stronger visual encoder, namely SigLIP(Zhai et al., 2023); a stronger LLM, namely Mistral-7b (Jiang et al., 2023); and a stronger diffuser, namely SDv1.4 (Rombach et al., 2022). L stands for Linear projection, and Q stands for Q-former resampler.
@@
In this paper, we present generative visual instruction tuning, an approach in which we teach a Large Multimodal Model (LMM) image understanding, image generation, and image editing tasks without diminishing the performance of each individual capability. (See Fig. 1 for an overview of our method.) To our knowledge, this is the first time such capability has been achieved, and our findings pave the way for building a general-purpose visual assistant. Our contributions are the following:
@@
  • Generative multimodal instruction-following data. Inspired by Liu et al. (2023b), which curated an instruction set for image understanding tasks, we curate a multimodal instruction tuning set that combines image understanding, image generation, and image editing data.
  • A single composite model, i.e. GenLLaVA, which unifies visual understanding and generation using open-source models. GenLLaVA is trained using a single-stage training recipe, unlike its predecessor LLaVA Liu et al. (2024a).
@@
  • Open source. We will publicly release our generated multimodal instruction data, code to replicate our results, model checkpoints, and a visual chat demo.
@@ 2 Related Work
Large Multimodal Models (LMMs).
Large Multimodal Models (LMMs) refer to large language models that can understand various modalities beyond human language. Some research efforts are focused on combining image, audio, video, and other modalities with language  (Zhan et al., 2024; Lu et al., 2024a), while others aim to enhance the fusion of vision knowledge and language. For example, BLIP-2 (Li et al., 2023) created a large-scale image captioning dataset and paired a language model with a vision encoder to produce a robust multimodal model. Following this, LLaVA (Liu et al., 2023b) developed a cost-effective approach to train an advanced LLM through visual instruction tuning. Although LLaVA-NeXT (Liu et al., 2024b) improved performance for single-image tasks, it required over 2,000 tokens per image, which is about four times more than the original LLaVA. More recent models such as QwenVL (Bai et al., 2023), CogVLM (Wang et al., 2023b), and Yi-VL (Young et al., 2024) follow architectures similar to those of LLaVA. Our proposed method not only focuses on models for multimodal understanding but also on adding generative capabilities to such models.
Diffusion-based LMMs for visual generation
We review works that combine diffusion with autoregressive prediction to create LMMs for generative tasks. For instance, GILL (Koh et al., 2024) translates the hidden representations of an LLM into embeddings that correspond to a text-to-image model by learning a neural network to perform efficient mapping using the text encoder of the diffusion model. MGIE (Fu et al., 2024) adapts the text embedder, image input adapter, and LM head output parameters of an LMM jointly with a diffusion model for image editing from instructions. DreamLLM (Dong et al., 2024) uses the same paradigm as GILL and MGIE but instead trains on interleaves documents for visual generation and understanding synergy. Transfusion (Zhou et al., 2024) combines the next-token production objective with the using a bidirectional casual mask on the visual tokens in a single unified model to understand and generate both discrete and continuous modalities. Show-O (Xie et al., 2024) is concurrent to Transfusion and similarly uses a bidirectional casual mask on the visual tokens but uses an extra masking loss similar to MaskGIT (Chang et al., 2022).
Token-based LMMs for visual generation
We review works that project visual features into discrete tokens and use next-token prediction for generative tasks. AnyGPT (Zhan et al., 2024) discretizes data from multiple modalities, extends the existing LLM vocabulary to add the extra modalities, and incorporates new randomly initialized parameters that enable additional input embeddings and prediction outputs. Codi-2 (Tang et al., 2024) follows a similar line and adds multiple decoders for image, audio, and video generation to condition an LLM to generate interleaved outputs from different modalities. CM3leon (Yu et al., 2023) proposes an early-fusion token-based decoder-only mixed modal model based on the CM3 architecture that is capable of both text and image generation and editing. Unified-IO 2 (Lu et al., 2024a), Chameleon (Chameleon Team, 2024) and GPT-4o OpenAI (2024) take the early-fusion fully multimodal approach training the model from scratch to be able to expand the number of supported tasks and modalities.
@@ 3 Method
Figure 2: Editing capabilities of our model. GPT4 currently uses a version of the DALLE text-to-image model as a tool and, hence, is not directly able to edit images. GPT4o instead uses tools through Python-generated code to accomplish the requested action. Our model, GenLLaVA, connects input features obtained from CLIP to a language model that also produces output embeddings for a text-to-image StableDiffusion model, achieving an end-to-end editing task with a multimodal model.
@@ 3.1 Background: Large Multimodal Models
Large language models (LLMs) excel in natural language generation, while Large Multimodal Models enhance LLMs with the ability to interpret images and respond accordingly. Built upon a pre-trained LLM, the LMM incorporates a visual encoder (e.g., CLIP (Radford et al., 2021)) to derive visual features $f$, along with an adapter $\mathcal{W}$ that maps $f$ into the language domain. Following the training methodology of LLaVA (Liu et al., 2023b), this process is encapsulated in the equation:
@@
where $l$ represents the number of tokens within $\mathcal{C}$. The set $\mathcal{C}$ can represent an image caption (Features Alignment) or multimodal instruction-following data (Instruction Tuning). The LMM employs the standard autoregressive method for next-token prediction, allowing it to function as a visual assistant across diverse tasks such as visual question answering and complex reasoning. Despite gaining visual perceptive abilities through this training, its responses are currently constrained to text.
@@ 3.2 Visual Generation in Large Multimodal Models
We append $N$ visual tokens [IMG] after the instruction $\mathcal{E}$, with their word embeddings being trainable. The LMM learns to generate these tokens through its language modeling (LM) head. These visual tokens represent visual-related instruction comprehension within $\mathcal{E}$ and form a bridge between the language and vision modalities. We follow the same visual generation framework of GILL (Koh et al., 2024) and MGIE (Fu et al., 2024) in extracting visual features, which we summarize here for succinctness.
@@
To guide image generation with the visual latent information $\mathcal{U}$, we employ a latent diffusion model (Rombach et al., 2022), incorporating a variational autoencoder (VAE) for handling denoising diffusion in the latent space. First, we encode the desired visual output via the diffusion model encoder $o=\text{Enc}_{\text{VAE}}(\mathcal{O})$; this output may be intended for image generation or editing tasks. The diffusion process progressively introduces noise into $o$ as $z_{t}$, increasing the noise level over timesteps $t$. We then train the UNet $\epsilon_{\theta}$ to predict the added noise (Ho et al., 2020). The diffusion process is conditioned on the visual latent information $\mathcal{U}$ through a cross-attention layer, defined as $\text{Attention}(Q,K,V)=\text{softmax}(\frac{QK^{T}}{\sqrt{\text{dim}}})\cdot V$, where:
$Q=W^{(i)}_{Q}\cdot\varphi_{i}(z_{t}),K=W^{(i)}_{K}\cdot\mathcal{U},V=W^{(i)}_{V}\cdot\mathcal{U},$ (3)
with $\varphi$ representing the flattening operation, and $W^{(i)}_{Q}$, $W^{(i)}_{K}$, and $W^{(i)}_{V}$ being learnable attention matrices. We apply classifier-free guidance (Ho & Salimans, 2021), where the score estimation $s_{\theta}$ is extrapolated to deviate from the unconditional $\varnothing$, following standard practices in diffusion models.
@@ 4.1 Generative Visual Instruction Data
We use the publicly available ShareGPT (ShareGPT, 2023) dataset, which was used to train the Vicuna LLM (Chiang et al., 2023). This dataset contains mostly English natural language conversations but also contains code and markdown. We filter inappropriate or low-quality entries following the same methodology of Chiang et al. (2023). As a final preprocessing step, entries that surpass 2048 tokens are truncated rather than split into multiple conversations. This results in $\sim$40K conversations.
Image Editing. [9.63%]
We create a subset of the Instruction Prompt-to-Prompt dataset (IPr2Pr) (Brooks et al., 2023) for editing our editing data. We use $\sim$200K from the CLIP-filtered data version of IPr2Pr, where editing instructions are generated by GPT-3, and images are synthesized by the Prompt-to-Prompt model (Hertz et al., 2023).
@@
For text-to-image generation, we use the same image & text pairs that were used to pre-train the LLaVA model. This dataset, named LLaVA-Pretrain, is inverted and presented to our model in the format (caption, image) with a dynamically pre-generated e.g.Please generate an image of caption”. These prefixes are created using the GPT4 language model. This dataset contains $\sim$558K data points originally sourced from the LAION (Schuhmann et al., 2022), SBU (Ordonez et al., 2011), and CC3M (Changpinyo et al., 2021) datasets and captioned by the BLIP-2 model (Li et al., 2023).
Image Understanding.[61.56%]
@@
For image understanding, we combine the dataset used to fine-tune the LLaVA model. This dataset, named LLaVA-Finetune, contains $\sim$665K samples. We also add the LVIS-INSTRUCT4V (Wang et al., 2023a) dataset—a new visual instruction tuning dataset constructed in the same way as the original LlaVA dataset but using GPT4-V (Achiam et al., 2023) as the captioner instead of BLIP-2 (Li et al., 2023). We remove duplicates from the resulting dataset; this results in $\sim$880K samples. We additionally add the following instruction datasets:
  • LRV-Instruction (Liu et al., 2023a) ($\sim$80K) a diagram undestanding and hallucination reduction dataset.
  • laion-gpt4v-dataset ($\sim$15K) a subset of the LAION (Schuhmann et al., 2022) dataset with high-quality captions created using GPT-4V (Achiam et al., 2023).
  • ShareGPT4V (Chen et al., 2024a) ($\sim$100K) a conversational dataset created using publicly available conversations that users had with the GPT-4V model.
  • Datasets for documents, chart and OCR understanding such as DocVQA (Mathew et al., 2021)($\sim$50K), SynDog-EN (Kim et al., 2022)($\sim$65K), ChartQA (Masry et al., 2022)($\sim$23K), DVQA (Kafle et al., 2018) ($\sim$50K) and AI2D (Kembhavi et al., 2016) ($\sim$15K).
4.2 Training details.
In this section, we evaluate our model on a broad range of tasks that require visual understanding and generation. We do not perform task-specific finetuning in any experiments. The Supplementary section details additional results on GenLLaVA ’s instruction capabilities.
We adopt LLaVA-v1.5-7B (Liu et al., 2024a) architecture, then tune it on the constructed GVIT-mix-2076K. We named this model GenLLaVA and it is made of the following components:
  • Image Processing & Visual Representations. We implement all image processing logic using the default image transforms provided by torchvision and the TIMM library (Wightman, 2019). We normalize pixel values using the default ImageNet values. The default backbone employed by all visual representations $\text{Enc}_{\text{vis}}$ that we evaluate in this work is a Vision Transformer (Dosovitskiy et al., 2021); we extract patch features from the penultimate layer, following LLaVA (Liu et al., 2023b).
  • Vision-Language Projector. We use a simple 2-layer GELU MLP as the projector $\mathcal{W}$, which projects each patch independently into the embedding space of the language model.
  • Language Model. We choose the Mistral-7B LLM (Jiang et al., 2023). In order to combine the projected visual patch embeddings, we perform simple sequence-wise concatenation, placing the patch embeddings before the text embeddings.
  • Visual Generation Head. The generation head $\mathcal{T}$ is a lightweight 4-layer encoder-decoder Transformer, which takes word embeddings $e$ and hidden states $h$ from the [IMG] tokens, as well as $L$ learnable query tokens as the input and generates the visual latent $\mathcal{U}$, we use the $L=77$, and the dimension of each $u_{t}\in\mathcal{U}$ is $768$.
  • Diffusion Image Decoder. We adopt Stable Diffusion v1.4 (SDv1.4) (Rombach et al., 2022) trained on 512×512 resolution. Similar to the visual encoder, the SD model is frozen without any modifications or training throughout the whole process.
Table 1: Main result. Comparison of various models across advanced knowledge and general understanding. MGIE was not originally designed for these tasks, as it is purely an editing model. For VQA, we take the generated caption as the answer, and when asking it to generate entirely new images, we provide a blank image as the prompt. We intend to show that models lose previous capabilities when we add a new one.
Model name Advanced Knowledge General Understanding Editing Generation
MathVista MMMU MMVet SEED-B MMB EVR CC3M COCO
GILL (Koh et al., 2024) 18.6 26.8 13.0 29.4 38.2 30.4 15.3 0.67
AnyGPT (Zhan et al., 2024) 24.4 24.0 14.8 28.0 36.0 40.3 14.3 0.65
MGIE (Fu et al., 2024) 15.5 25.6 13.0 28.8 6.6 71.5 13.6 0.66
Chameleon (Chameleon Team, 2024) 22.3 22.4 10.3 30.5 15.4 - 10.2 0.78
Unified-IO 2 (Lu et al., 2024a) 28.3 35.5 36.6 61.6 57.9 50.2 13.4 0.72
GenLLaVA (Ours) 30.5 37.1 35.8 64.5 66.8 66.9 12.5 0.73
We implement our training codebase in PyTorch. We train all models in BF16 mixed precision. For a fair comparison, the rest of the model training protocol is kept unchanged from the original LLaVA. Generative Visual Instruction tuning takes about 48 hours for both full-parameter tuning and LoRA tuning on 8 NVIDIA Tesla A100 GPUs, each with 48GB memory, with DeepSpeed ZeRO Stage 3 (Rajbhandari et al., 2020) to distribute training across GPUs.
Single-stage training.
Unlike its predecessor LLaVA (Liu et al., 2023b), our model does not use a two-stage training pipeline and instead directly finetunes the Vision-Language projector, the Language model, and the Visual Generation Head. We found that the logical extension of the original pipeline—a three-stage training pipeline—consisting of (1) multimodal alignment, (2) instruction tuning, and (3) image generation tuning to teach the models progressively simply does not work, and the model performance on visual understanding tasks decreases significantly. We instead choose a single-stage pipeline as it has been shown to work previously in the work of Karamcheti et al. (2024). This comes with other unintentional advantages: training cost is reduced by 20%, and we can use the LLaVA-Pretrain data as image-to-text data instead of having to collect more.
Task Tokens.
To perform unified learning on multimodal understanding and generation, we introduce special tokens, which we name task tokens, to format the data. Specifically, we create the tokens [T2I] and [I2T] to indicate the learning (generation or understanding) task for the input sequence. We keep the original special token from the language model that indicates the start and end of the text. Similarly, [SOI] and [EOI] are pre-defined special tokens marking the start and end of visual tokens for generation. Without these task tokens, the model has trouble inferring the user intention and would generate an image when it is not necessary. We remark that this methodology is not new and has been used before by others (Lu et al., 2024a; Xie et al., 2024).
4.3 Evaluation Details.
Visual Understanding
We evaluate vision-language performance and compare it against other generalist models, i.e., models that can do visual generation and understating. Results on a collection of 5 vision/language benchmarks are shown in Table 1. These benchmarks are designed to probe advanced knowledge and general understanding. MMBench (Liu et al., 2024c) evaluates answer robustness through comprehensive shuffling of multiple-choice options. SEED-Bench (Li et al., 2024) tests model performance on images and videos with multiple-choice questions. MM-Vet (Yu et al., 2024) evaluates the ability to engage in visual conversations and assess the accuracy and helpfulness of responses. Mathvista (Lu et al., 2024b) consolidates mathematical reasoning benchmarks, focusing on logical and algebraic reasoning with puzzle tests. MMMU (Yue et al., 2024) covers 57 subjects across STEM, humanities, and social sciences, ranging from elementary to advanced professional levels, and tests both world knowledge and problem-solving skills.
@@
For visual generation evaluation, we measured the Fréchet Inception Distance (FID) (Heusel et al., 2017) on the CC3M validation set (Changpinyo et al., 2021) (which is a measure of image realism) and CLIP Similarity on the MS-COCO dataset (Lin et al., 2014) (which is a measure of the alignment between the text prompt and the generated image). For image editing, we measured the DINOScore on the 5.7K validation examples of the EVR dataset (Tan et al., 2019) following the same protocol as Fu et al. (2024).